3 identical point charges, each charge coulomb are placed at the vertical of an equilateral triangle 10cm apart. Calculate the force on each charge.
1
Expert's answer
2019-01-22T11:11:06-0500

q=1 C.

a=10 cm=0.1 m

F_12=F_13=F=(kq^2)/a^2

F_1x=F+F cos⁡60=F+1/2 F=3/2 F

F_1y=F sin⁡60=√3/2 F

Thus,

F_1=√((3/2 F)^2+(√3/2 F)^2 )=√3 F

F_1=√3 (kq^2)/a^2

F_1=√3 (9∙〖10〗^9 ) (1/0.1)^2=1.56∙〖10〗^12 N.